If a lossless transformer has 1000 turns for a primary winding and 100 turns for the secondary winding, and the voltage at the secondary terminal is 10 volts, what is the voltage at the primary terminal?
Answer #1

Transformation ratio = Ns/Np = Vs/Vp

100/1000 = 10/Vp

Voltage of primary terminal, Vp = 100

Comment in case any doubt please rate my answer....